OpenStudy (anonymous):

As you can see, the biker and his friend are having a hard time getting up the hill. They would like to get to the top safely. Two ordered pairs on the hill are given as (6, 3) and (8, 7). Part 1: Find the slope of the hill.

OpenStudy (anonymous):

Part 2: Now find the equation using the point slope formula. Point Slope Formula: y – y1 = m (x – x1) Part 3: Give an example of an equation for a line parallel to your answer from Part 2. Part 4: Give an example of an equation for a line perpendicular to your answer from Part 2.

OpenStudy (aripotta):

give me a couple minutes, please. too many people are asking for my help >.<

OpenStudy (aripotta):

ok, have you found the slope yet?

OpenStudy (anonymous):

no.. :(

OpenStudy (aripotta):

you know how to, right?

OpenStudy (aripotta):



OpenStudy (anonymous):

it is 2

OpenStudy (aripotta):

mhmm

OpenStudy (anonymous):

so now to part 2

OpenStudy (aripotta):

so, which coordinate do you wanna work with? the first or second?

OpenStudy (anonymous):

first

OpenStudy (aripotta):

ok. so the point-slope formula will be: y - 3 = 2(x - 6)

OpenStudy (aripotta):

do you understand that?

OpenStudy (anonymous):

yes. so what do i do now

OpenStudy (aripotta):

ummm

OpenStudy (aripotta):

we need to come up with 2 points that would have the same slope as 2

OpenStudy (aripotta):

so...how about (6,5) and (8,9)? would they work?

OpenStudy (anonymous):

sure

OpenStudy (aripotta):

ok so their equation would be: y - 5 = 2(x - 6)

OpenStudy (aripotta):

i'm not sure how to do the last part though :/

OpenStudy (anonymous):

wait were on part 2 ?

OpenStudy (aripotta):

lol i just did parts 1-3

OpenStudy (anonymous):

so the 2nd part is cause im confused

OpenStudy (anonymous):

: y - 5 = 2(x - 6) part 2 ?

OpenStudy (aripotta):

part 1: 2 part 2: y - 3 = 2(x - 6) part 3: y - 5 = 2(x - 6)